# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ Appendix B CLEAR Algorithm Details
Appendix C Experimental Setup: Details
C.1 Additional Baselines
The following baselines provide further context and comparisons, omitted from the main paper for brevity. All compared methods (except UACQR) use the same median $\hat{f}(x)=\hat{f}_{\text{PCS}}(x)$ obtained from PCS. (Obviously, all the results from Section E.1 use PCS variant (a), from Section E.2 use PCS variant (b), from Section E.3 use PCS variant (c).) This allows us to isolate the effect of different uncertainty quantification methods.
@@ C.1 Additional Baselines
    UACQR-S and UACQR-P For variant (c), instead of computing $C_{\gamma_{1}=1}(x)$ and $C_{\lambda=1}(x)$ with CLEAR, we directly use the implementation from [19] to assess our performance against this alternative. Since this is only relevant for variant (c), we use the exact same configuration as the aleatoric part for QRF (see the following subsection and Table 3).
C.2 PCS Implementation Details
@@
The point predictor $\hat{f}(x)$ and the raw epistemic uncertainty contributions $\hat{q}^{\text{epi}}_{\alpha/2}(x)$ and $\hat{q}^{\text{epi}}_{1-\alpha/2}(x)$ supplied to the CLEAR method are derived from this final ensemble of 100 model instances. As detailed in Section 2.2, $\hat{f}(x)$ is the pointwise empirical median of the ensemble’s predictions. The epistemic uncertainty terms represent the pointwise distances from this median to the ensemble’s empirical $\alpha/2$ and $21-\alpha/$ 2 quantiles, respectively.
@@
The standalone PCS baseline method involves a distinct calibration process. From the ensemble’s raw pointwise $\alpha/2$ and $21-\alpha/$ 2 quantile predictions ($\hat{f}_{\alpha/2}(x)$ and $\hat{f}_{1-\alpha/2}(x)$), a single, global multiplicative calibration factor, $\gamma_{\text{PCS}}$, is computed. This $\gamma_{\text{PCS}}$ is the smallest value ensuring that prediction intervals, formed by scaling the raw epistemic uncertainty around $\hat{f}(x)$ (i.e., $[\hat{f}(x)-\gamma_{\text{PCS}}(\hat{f}(x)-\hat{f}_{\alpha/2}(x)),\,\hat{f}(x)+\gamma_{\text{PCS}}(\hat{f}_{1-\alpha/2}(x)-\hat{f}(x))]$), achieve the target $1-\alpha$ coverage on the validation set, incorporating the standard finite-sample correction. This $\gamma_{\text{PCS}}$ is then applied to generate the PCS baseline intervals on the test set. It is important to reiterate that this $\gamma_{\text{PCS}}$ is separate and computed independently from $\gamma_{1}$ and $\lambda$ parameters optimized within the CLEAR framework.
C.3 Metrics
This appendix provides detailed definitions for the evaluation metrics used in the main paper. We consider test data ${(X_{i},Y_{i})}_{i=1}^{N}$ and prediction intervals $[L_{i},U_{i}]$.
@@ C.3 Metrics
  • Quantile Loss (also known as pinball loss): Evaluates the accuracy of predicted quantiles by penalizing both under- and overestimation. It reflects a trade-off between coverage (PICP) and interval width (NIW), rewarding narrow intervals that still maintain proper coverage.
@@
    $QuantileLoss(L,U)=\sum_{i=1}^{N}\left[QL_{\alpha/2}(Y_{i},L(X_{i}))+QL_{1-\alpha/2}(Y_{i},U(X_{i}))\right]/2.$
  • Average Interval Score Loss (AISL) [61]: This score balances interval width with coverage penalties, defined as
[…]
@@
Appendix D Simulations
D.1 Heteroskedastic Case
In  Section 3.1, we focused on the univariate homoskedastic setting ($d=1$ and $\sigma(x)=1$). Here, we briefly report results for heteroskedastic data, which show similar patterns. Figures 4 and 5 illustrate the conditional coverage and interval width of the algorithms under two heteroskedastic noise structures: $\sigma_{2}(x)=1+|x|$ and $\sigma_{3}(x)=1+\frac{1}{1+x^{2}}$. As mentioned, the results are are analogous to the homoskedastic setting in  Section 3.1.
D.2 Multivariate case
D.2.1 Test point generation
D.2.2 Results
Appendix E Full results on real-world datasets

### v2

@@ B.5 CLEAR Algorithm: Details
Appendix C Simulations: Details
C.1 Heteroskedastic Case
In  Section 3.1, we focused on the univariate homoskedastic setting ($d=1$ and $\sigma(x)=1$). Here, we briefly report results for heteroskedastic data, which show similar patterns. Figures 4 and 5 illustrate the conditional coverage and interval width of the algorithms under two heteroskedastic noise structures: $\sigma_{2}(x)=1+|x|$ and $\sigma_{3}(x)=1+\frac{1}{1+x^{2}}$. As mentioned, the results are analogous to the homoskedastic setting in  Section 3.1.
C.2 Multivariate Case
C.2.1 Test point generation
C.2.2 Results
Appendix D Experimental Setup: Details
D.1 Real-World Data
To evaluate our method in real-world scenarios, we apply it to 17 publicly available regression datasets curated by [30] that includes various domains such as housing, energy, materials science, and healthcare [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. Description of the datasets can be found in Table 11.
D.2 Baselines for Variants (a), (b), (c)
The following baselines provide further context and comparisons, omitted from the main paper for brevity. All compared methods (except UACQR) use the same median $\hat{f}(x)=\hat{f}_{\text{PCS}}(x)$ obtained from PCS. The results from Section F.1 use PCS variant (a), Section F.2 uses PCS variant (b), and Section F.3 uses PCS variant (c). This allows us to isolate the effect of different UQ methods.
@@ D.2 Baselines for Variants (a), (b), (c)
    UACQR-S and UACQR-P For variant (c), instead of computing $C_{\gamma_{1}=1}(x)$ and $C_{\lambda=1}(x)$ with CLEAR, we directly use the implementation from [19] to assess our performance against this alternative. Since this is only relevant for variant (c), we use the exact same configuration as the aleatoric part for QRF (see Section D.4 and Table 4).
D.3 Deep Ensembles and Simultaneous Quantile Regression Implementation
To demonstrate CLEAR’s versatility beyond tree-based methods, we implement state-of-the-art deep learning approaches for uncertainty quantification. These neural network-based methods offer complementary strengths to our primary PCS framework, particularly in capturing complex nonlinear relationships and heteroskedastic noise patterns.
D.3.1 Deep Ensembles for Epistemic Uncertainty
Our deep ensemble implementation follows [22] with several enhancements for improved diversity and calibration. Each ensemble consists of $M=5$ neural networks (adaptive based on dataset size), with the following specifications:
  • Architecture: Each network employs a fully-connected architecture with hidden layers of sizes (256, 128), ReLU activations, batch normalization, and dropout (rate=0.1). Skip connections are incorporated to stabilize training.
  • Diversity strategies: To promote ensemble diversity, we employ: (i) different random initializations, (ii) bootstrap sampling where each member trains on different subsamples, (iii) varied learning rates, (iv) different learning rate schedules, and (v) small input noise augmentation.
  • Training: Each network is trained for up to 1500 epochs using Adam optimizer with learning rate $10^{-3}$, early stopping (patience=50), weight decay ($10^{-5}$), and batch size 64.
  • Calibration: Following training, we apply a multiplicative calibration factor computed on the validation set using a grid search over c-values to ensure the ensemble achieves target coverage.
The epistemic uncertainty is quantified through the empirical quantiles of the ensemble predictions: $\hat{q}^{\text{epi}}_{\alpha/2}(x)=\hat{f}(x)-Q_{\alpha/2}(\{\hat{f}_{m}(x)\}_{m=1}^{M})$ and $\hat{q}^{\text{epi}}_{1-\alpha/2}(x)=Q_{1-\alpha/2}(\{\hat{f}_{m}(x)\}_{m=1}^{M})-\hat{f}(x)$, where $\hat{f}(x)$ is the ensemble median and $Q_{\tau}$ denotes the empirical quantile function.
D.3.2 Simultaneous Quantile Regression for Aleatoric Uncertainty
For aleatoric uncertainty, we implement simultaneous quantile regression (SQR) following [13], which directly models multiple conditional quantiles through a single neural network with specialized architecture:
  • Architecture: The network uses hidden layers of sizes (256, 256, 128) with LeakyReLU activations (negative slope=0.01), layer normalization, and dropout (rate=0.2). The output layer produces three values corresponding to the $\alpha/2$, 0.5, and $1-\alpha/2$ quantiles.
  • Loss function: We minimize a combined loss consisting of: (i) the pinball loss for each quantile, and (ii) a crossing penalty term that encourages monotonicity of quantiles: $\mathcal{L}_{\text{crossing}}=\text{ReLU}(q_{\alpha/2}-q_{0.5})+\text{ReLU}(q_{0.5}-q_{1-\alpha/2})$.
  • Training: The model is trained for up to 3000 epochs using Adam optimizer with learning rate $5\times 10^{-4}$, cosine annealing schedule, and early stopping (patience=200). Gradient clipping (max norm=1.0) prevents training instabilities.
  • Ensemble averaging: To improve stability, we train an ensemble of SQR models with different random seeds and average their quantile predictions.
The aleatoric uncertainty estimates are computed as $\hat{q}^{\text{ale}}_{\alpha/2}(x)=q_{0.5}(x)-q_{\alpha/2}(x)$ and $\hat{q}^{\text{ale}}_{1-\alpha/2}(x)=q_{1-\alpha/2}(x)-q_{0.5}(x)$, ensuring consistency with the median prediction.
D.3.3 Integration with CLEAR
The DE and SQR components are integrated into CLEAR using the same calibration procedure as our PCS-UQ (model selection and tree-based methods). The epistemic estimates from DE and aleatoric estimates from SQR are combined using the optimized parameters $\lambda$ and $\gamma_{1}$ according to Equation (3). The conformal calibration step ensures marginal coverage while the adaptive $\lambda$ selection balances the relative contributions of epistemic and aleatoric uncertainties, accounting for potential scale differences between neural network and tree-based estimators.
D.4 PCS Implementation: Details
@@
The point predictor $\hat{f}(x)$ and the raw epistemic uncertainty contributions $\hat{q}^{\text{epi}}_{\alpha/2}(x)$ and $\hat{q}^{\text{epi}}_{1-\alpha/2}(x)$ supplied to the CLEAR method are derived from this final ensemble of 100 model instances. As detailed in Section 2.2, $\hat{f}(x)$ is the pointwise empirical median of the ensemble’s predictions. The epistemic uncertainty terms represent the pointwise distances from this median to the ensemble’s empirical $\alpha/2$ and $1-\alpha/2$ quantiles, respectively.
@@
The standalone PCS baseline method involves a distinct calibration process. From the ensemble’s raw pointwise $\alpha/2$ and $1-\alpha/2$ quantile predictions ($\hat{f}_{\alpha/2}(x)$ and $\hat{f}_{1-\alpha/2}(x)$), a single, global multiplicative calibration factor, $\gamma_{\text{PCS}}$, is computed. This $\gamma_{\text{PCS}}$ is the smallest value ensuring that prediction intervals, formed by scaling the raw epistemic uncertainty around $\hat{f}(x)$ (i.e., $\hat{f}(x)-\gamma_{\text{PCS}}(\hat{f}(x)-\hat{f}_{\alpha/2}(x)),\,\hat{f}(x)+\gamma_{\text{PCS}}(\hat{f}_{1-\alpha/2}(x)-\hat{f}(x))]$), achieve the target $1-\alpha$ coverage on the validation set, incorporating the standard finite-sample correction. This $\gamma_{\text{PCS}}$ is then applied to generate the PCS baseline intervals on the test set. It is important to reiterate that this $\gamma_{\text{PCS}}$ is separate and computed independently from $\gamma_{1}$ and $\lambda$ parameters optimized within the CLEAR framework.
D.5 Metrics
This appendix provides detailed definitions for the evaluation metrics used in the main paper. We consider test data ${(X_{i},Y_{i})}_{i=1}^{N}$ and prediction intervals $L_{i},U_{i}]$.
@@ D.5 Metrics
  • Quantile Loss (also known as pinball loss): Evaluates the accuracy of predicted quantiles by penalizing both under- and overestimation. It reflects a trade-off between coverage (PICP) and interval width (NIW), rewarding narrow intervals that still maintain proper coverage and penalizing data points that are far outside the intervals (see Section B.3 for more details).
@@
    $\text{QuantileLoss}(L,U)=\sum_{i=1}^{N}\left[QL_{\alpha/2}(Y_{i},L(X_{i}))+QL_{1-\alpha/2}(Y_{i},U(X_{i}))\right]/2.$
  • Average Interval Score Loss (AISL) [110]: This score balances interval width with coverage penalties (with similar properties as the quantile loss explained in Section B.3), defined as
    $\text{AISL}(L,U)=\frac{1}{N}\sum_{i=1}^{N}\bigg[$ $(U_{i}-L_{i})+\frac{2}{\alpha}(L_{i}-Y_{i})\,\mathds{1}\{Y_{i}<L_{i}\}+\frac{2}{\alpha}(Y_{i}-U_{i})\,\mathds{1}\{Y_{i}>U_{i}\}\bigg],$
@@
Appendix E CLEAR with DE and SQR: Results on Real-World Data
This section presents the comprehensive experimental results of CLEAR when combined with deep learning-based uncertainty estimators: deep ensembles (DE) for epistemic uncertainty and simultaneous quantile regression (SQR) for aleatoric uncertainty. The experiments further validate CLEAR’s generality beyond the PCS and CQR methods that are presented in the body of our paper. The results demonstrate that CLEAR remains effective across different modeling paradigms. We evaluate both the neural baselines individually and their integration through CLEAR, comparing against conformalized versions to assess the added value of our dual-parameter calibration approach. Results are reported across all 17 datasets with 95% nominal coverage, using up to 10 random seeds for robustness.
Table 6: DE-SQR PICP at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.94 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
airfoil 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.98 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.91 $\pm$ 0.01 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.94 $\pm$ 0.02 0.95 $\pm$ 0.02 0.98 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.03
elevator 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.94 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
energy_efficiency 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.99 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.03
insurance 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02
kin8nm 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.98 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
naval_propulsion 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
qsar 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.93 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
Table 7: DE-SQR NIW at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 0.201 $\pm$ 0.017 0.284 $\pm$ 0.025 0.196 $\pm$ 0.014 0.285 $\pm$ 0.025 0.203 $\pm$ 0.015
airfoil 0.185 $\pm$ 0.023 0.223 $\pm$ 0.023 0.349 $\pm$ 0.025 0.226 $\pm$ 0.023 0.311 $\pm$ 0.025
allstate 0.333 $\pm$ 0.061 0.380 $\pm$ 0.073 0.283 $\pm$ 0.043 0.383 $\pm$ 0.075 0.324 $\pm$ 0.058
ca_housing 0.357 $\pm$ 9.1e-03 0.502 $\pm$ 0.020 0.378 $\pm$ 8.0e-03 0.502 $\pm$ 0.020 0.374 $\pm$ 4.0e-03
computer 0.087 $\pm$ 3.2e-03 0.121 $\pm$ 6.3e-03 0.104 $\pm$ 8.9e-03 0.122 $\pm$ 6.1e-03 0.106 $\pm$ 5.7e-03
concrete 0.264 $\pm$ 0.033 0.365 $\pm$ 0.068 0.441 $\pm$ 0.041 0.369 $\pm$ 0.067 0.392 $\pm$ 0.069
elevator 0.117 $\pm$ 6.6e-03 0.174 $\pm$ 0.011 0.114 $\pm$ 0.011 0.174 $\pm$ 0.011 0.117 $\pm$ 8.8e-03
energy_efficiency 0.085 $\pm$ 0.012 0.115 $\pm$ 0.020 0.480 $\pm$ 0.059 0.120 $\pm$ 0.020 0.368 $\pm$ 0.036
insurance 0.393 $\pm$ 0.044 0.523 $\pm$ 0.135 0.397 $\pm$ 0.073 0.558 $\pm$ 0.141 0.406 $\pm$ 0.085
kin8nm 0.197 $\pm$ 8.2e-03 0.260 $\pm$ 9.5e-03 0.246 $\pm$ 8.7e-03 0.261 $\pm$ 9.6e-03 0.208 $\pm$ 8.8e-03
miami_housing 0.093 $\pm$ 2.9e-03 0.131 $\pm$ 5.6e-03 0.101 $\pm$ 2.8e-03 0.131 $\pm$ 5.7e-03 0.100 $\pm$ 1.8e-03
naval_propulsion 1.5e-03 $\pm$ 4.2e-04 5.4e-03 $\pm$ 1.1e-03 3.7e-03 $\pm$ 1.3e-04 5.4e-03 $\pm$ 1.1e-03 1.7e-03 $\pm$ 1.5e-04
parkinsons 0.378 $\pm$ 0.014 0.476 $\pm$ 0.027 0.419 $\pm$ 0.021 0.479 $\pm$ 0.027 0.403 $\pm$ 0.017
powerplant 0.186 $\pm$ 6.7e-03 0.266 $\pm$ 0.012 0.204 $\pm$ 6.8e-03 0.267 $\pm$ 0.013 0.202 $\pm$ 7.3e-03
qsar 0.410 $\pm$ 0.136 0.508 $\pm$ 0.178 0.402 $\pm$ 0.135 0.509 $\pm$ 0.179 0.456 $\pm$ 0.153
sulfur 0.105 $\pm$ 8.5e-03 0.199 $\pm$ 0.015 0.120 $\pm$ 7.4e-03 0.200 $\pm$ 0.015 0.120 $\pm$ 8.3e-03
superconductor 0.208 $\pm$ 0.024 0.295 $\pm$ 0.038 0.247 $\pm$ 0.027 0.295 $\pm$ 0.038 0.245 $\pm$ 0.028
Table 8: DE-SQR Quantile Loss at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 9.6e-06 $\pm$ 3.1e-07 1.3e-05 $\pm$ 3.5e-07 9.6e-06 $\pm$ 4.0e-07 1.3e-05 $\pm$ 3.5e-07 9.5e-06 $\pm$ 3.7e-07
airfoil 0.097 $\pm$ 6.8e-03 0.117 $\pm$ 0.010 0.162 $\pm$ 0.017 0.117 $\pm$ 0.010 0.153 $\pm$ 0.018
allstate 1.4e+02 $\pm$ 9.329 1.5e+02 $\pm$ 8.112 1.5e+02 $\pm$ 9.210 1.5e+02 $\pm$ 8.106 1.4e+02 $\pm$ 8.368
ca_housing 3.0e+03 $\pm$ 63.719 3.9e+03 $\pm$ 1.2e+02 3.0e+03 $\pm$ 54.590 3.9e+03 $\pm$ 1.2e+02 3.0e+03 $\pm$ 53.118
computer 0.146 $\pm$ 5.5e-03 0.183 $\pm$ 6.8e-03 0.162 $\pm$ 8.3e-03 0.183 $\pm$ 6.8e-03 0.162 $\pm$ 7.9e-03
concrete 0.348 $\pm$ 0.049 0.427 $\pm$ 0.056 0.467 $\pm$ 0.031 0.428 $\pm$ 0.057 0.458 $\pm$ 0.051
elevator 1.2e-04 $\pm$ 1.9e-06 1.6e-04 $\pm$ 3.5e-06 1.1e-04 $\pm$ 2.5e-06 1.6e-04 $\pm$ 3.5e-06 1.1e-04 $\pm$ 2.4e-06
energy_efficiency 0.052 $\pm$ 9.5e-03 0.067 $\pm$ 0.012 0.221 $\pm$ 0.023 0.067 $\pm$ 0.012 0.188 $\pm$ 0.020
insurance 3.9e+02 $\pm$ 42.155 5.3e+02 $\pm$ 97.010 3.9e+02 $\pm$ 37.867 5.4e+02 $\pm$ 93.793 3.9e+02 $\pm$ 45.372
kin8nm 4.1e-03 $\pm$ 4.4e-05 5.2e-03 $\pm$ 8.2e-05 4.5e-03 $\pm$ 5.7e-05 5.2e-03 $\pm$ 8.2e-05 4.2e-03 $\pm$ 7.1e-05
miami_housing 4.5e+03 $\pm$ 2.1e+02 5.4e+03 $\pm$ 1.7e+02 4.4e+03 $\pm$ 1.2e+02 5.4e+03 $\pm$ 1.7e+02 4.4e+03 $\pm$ 1.2e+02
naval_propulsion 3.6e-05 $\pm$ 9.4e-06 1.3e-04 $\pm$ 2.7e-05 8.2e-05 $\pm$ 2.9e-06 1.3e-04 $\pm$ 2.7e-05 4.4e-05 $\pm$ 4.3e-06
parkinsons 0.298 $\pm$ 0.014 0.343 $\pm$ 9.8e-03 0.291 $\pm$ 5.7e-03 0.343 $\pm$ 9.8e-03 0.289 $\pm$ 5.8e-03
powerplant 0.227 $\pm$ 0.013 0.301 $\pm$ 0.014 0.234 $\pm$ 0.012 0.301 $\pm$ 0.014 0.234 $\pm$ 0.013
qsar 0.055 $\pm$ 3.2e-03 0.064 $\pm$ 3.5e-03 0.061 $\pm$ 3.7e-03 0.064 $\pm$ 3.5e-03 0.060 $\pm$ 3.6e-03
sulfur 1.7e-03 $\pm$ 1.0e-04 2.7e-03 $\pm$ 1.9e-04 1.8e-03 $\pm$ 1.3e-04 2.7e-03 $\pm$ 1.9e-04 1.8e-03 $\pm$ 1.3e-04
superconductor 0.545 $\pm$ 0.024 0.707 $\pm$ 0.024 0.560 $\pm$ 0.017 0.707 $\pm$ 0.024 0.559 $\pm$ 0.018
Table 9: DE-SQR NCIW at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 0.205 $\pm$ 0.017 0.292 $\pm$ 0.027 0.205 $\pm$ 0.017 0.291 $\pm$ 0.026 0.204 $\pm$ 0.017
airfoil 0.173 $\pm$ 0.016 0.220 $\pm$ 0.025 0.290 $\pm$ 0.026 0.217 $\pm$ 0.025 0.293 $\pm$ 0.026
allstate 0.328 $\pm$ 0.048 0.384 $\pm$ 0.062 0.342 $\pm$ 0.051 0.381 $\pm$ 0.060 0.330 $\pm$ 0.045
ca_housing 0.354 $\pm$ 8.1e-03 0.501 $\pm$ 0.021 0.372 $\pm$ 7.1e-03 0.501 $\pm$ 0.021 0.373 $\pm$ 7.8e-03
computer 0.088 $\pm$ 1.7e-03 0.120 $\pm$ 5.0e-03 0.106 $\pm$ 6.5e-03 0.120 $\pm$ 5.1e-03 0.106 $\pm$ 7.1e-03
concrete 0.274 $\pm$ 0.026 0.345 $\pm$ 0.053 0.371 $\pm$ 0.036 0.339 $\pm$ 0.051 0.376 $\pm$ 0.040
elevator 0.117 $\pm$ 7.0e-03 0.172 $\pm$ 0.010 0.117 $\pm$ 9.0e-03 0.172 $\pm$ 0.010 0.116 $\pm$ 9.4e-03
energy_efficiency 0.079 $\pm$ 7.4e-03 0.114 $\pm$ 0.017 0.351 $\pm$ 0.049 0.106 $\pm$ 0.014 0.346 $\pm$ 0.046
insurance 0.364 $\pm$ 0.060 0.465 $\pm$ 0.125 0.351 $\pm$ 0.064 0.460 $\pm$ 0.122 0.350 $\pm$ 0.061
kin8nm 0.194 $\pm$ 6.6e-03 0.255 $\pm$ 5.6e-03 0.205 $\pm$ 6.3e-03 0.254 $\pm$ 5.3e-03 0.208 $\pm$ 6.3e-03
miami_housing 0.092 $\pm$ 3.5e-03 0.129 $\pm$ 7.8e-03 0.100 $\pm$ 2.6e-03 0.129 $\pm$ 7.7e-03 0.100 $\pm$ 2.9e-03
naval_propulsion 1.5e-03 $\pm$ 4.2e-04 5.4e-03 $\pm$ 1.2e-03 1.7e-03 $\pm$ 2.0e-04 5.4e-03 $\pm$ 1.2e-03 1.6e-03 $\pm$ 1.7e-04
parkinsons 0.382 $\pm$ 0.015 0.473 $\pm$ 0.025 0.398 $\pm$ 0.013 0.470 $\pm$ 0.022 0.398 $\pm$ 0.013
powerplant 0.190 $\pm$ 6.0e-03 0.272 $\pm$ 9.5e-03 0.202 $\pm$ 8.1e-03 0.272 $\pm$ 9.5e-03 0.202 $\pm$ 8.1e-03
qsar 0.407 $\pm$ 0.134 0.477 $\pm$ 0.150 0.450 $\pm$ 0.151 0.476 $\pm$ 0.150 0.441 $\pm$ 0.146
sulfur 0.104 $\pm$ 6.2e-03 0.198 $\pm$ 0.016 0.120 $\pm$ 7.1e-03 0.197 $\pm$ 0.015 0.120 $\pm$ 7.2e-03
superconductor 0.209 $\pm$ 0.024 0.303 $\pm$ 0.036 0.242 $\pm$ 0.027 0.302 $\pm$ 0.036 0.245 $\pm$ 0.027
Table 10: DE-SQR Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 7.7e-04 $\pm$ 2.5e-05 1.0e-03 $\pm$ 2.8e-05 7.7e-04 $\pm$ 3.2e-05 1.0e-03 $\pm$ 2.8e-05 7.6e-04 $\pm$ 2.9e-05
airfoil 7.761 $\pm$ 0.546 9.368 $\pm$ 0.826 12.937 $\pm$ 1.386 9.370 $\pm$ 0.813 12.279 $\pm$ 1.418
allstate 1.1e+04 $\pm$ 7.5e+02 1.2e+04 $\pm$ 6.5e+02 1.2e+04 $\pm$ 7.4e+02 1.2e+04 $\pm$ 6.5e+02 1.2e+04 $\pm$ 6.7e+02
ca_housing 2.4e+05 $\pm$ 5.1e+03 3.1e+05 $\pm$ 9.2e+03 2.4e+05 $\pm$ 4.4e+03 3.1e+05 $\pm$ 9.2e+03 2.4e+05 $\pm$ 4.2e+03
computer 11.669 $\pm$ 0.436 14.633 $\pm$ 0.543 12.938 $\pm$ 0.662 14.634 $\pm$ 0.545 12.926 $\pm$ 0.634
concrete 27.850 $\pm$ 3.946 34.179 $\pm$ 4.507 37.343 $\pm$ 2.464 34.203 $\pm$ 4.564 36.678 $\pm$ 4.051
elevator 9.2e-03 $\pm$ 1.5e-04 0.013 $\pm$ 2.8e-04 9.0e-03 $\pm$ 2.0e-04 0.013 $\pm$ 2.8e-04 9.0e-03 $\pm$ 1.9e-04
energy_efficiency 4.154 $\pm$ 0.758 5.369 $\pm$ 0.965 17.719 $\pm$ 1.871 5.387 $\pm$ 0.974 15.020 $\pm$ 1.595
insurance 3.1e+04 $\pm$ 3.4e+03 4.2e+04 $\pm$ 7.8e+03 3.1e+04 $\pm$ 3.0e+03 4.3e+04 $\pm$ 7.5e+03 3.1e+04 $\pm$ 3.6e+03
kin8nm 0.327 $\pm$ 3.5e-03 0.417 $\pm$ 6.5e-03 0.358 $\pm$ 4.6e-03 0.417 $\pm$ 6.6e-03 0.340 $\pm$ 5.7e-03
miami_housing 3.6e+05 $\pm$ 1.7e+04 4.3e+05 $\pm$ 1.4e+04 3.5e+05 $\pm$ 9.5e+03 4.3e+05 $\pm$ 1.4e+04 3.5e+05 $\pm$ 9.5e+03
naval_propulsion 2.9e-03 $\pm$ 7.5e-04 0.011 $\pm$ 2.2e-03 6.5e-03 $\pm$ 2.3e-04 0.011 $\pm$ 2.2e-03 3.5e-03 $\pm$ 3.5e-04
parkinsons 23.860 $\pm$ 1.124 27.437 $\pm$ 0.787 23.297 $\pm$ 0.456 27.449 $\pm$ 0.784 23.159 $\pm$ 0.461
powerplant 18.145 $\pm$ 1.005 24.110 $\pm$ 1.141 18.714 $\pm$ 0.990 24.106 $\pm$ 1.140 18.706 $\pm$ 1.002
qsar 4.419 $\pm$ 0.255 5.152 $\pm$ 0.278 4.875 $\pm$ 0.294 5.154 $\pm$ 0.278 4.800 $\pm$ 0.286
sulfur 0.134 $\pm$ 8.1e-03 0.213 $\pm$ 0.015 0.144 $\pm$ 0.010 0.213 $\pm$ 0.015 0.144 $\pm$ 0.010
superconductor 43.610 $\pm$ 1.937 56.594 $\pm$ 1.943 44.763 $\pm$ 1.388 56.591 $\pm$ 1.943 44.731 $\pm$ 1.406
Appendix F CLEAR with PCS and CQR: Results on Real-World Data